Response Questions Part A I& B Measuring tree height is not an easy task. How well does trunk diameter predict tree trees in an old-growth forest in Canada provides us with the information: The mean tree height is 15.6 meters (m) with standard deviation 13.4m; the mean diameter, measured at breastheight (1.3m above ground), is 23.4 centimeters (cm), and the standard deviation is 23.5 cm. The correlation between the height and diameter is very high:r-0.96 A survey of 958 live (A) Whatis the slope of the regression line to predict tree height (in meters) from trunk diameter (in centimeters)? (Enter youranswer rounded to four decimal places.) What is the intercept of the regression line to predict tree height (in meters) from trunk diameter (in centimeters)? (Enter youranswer rounded to three decimal places Draw a graph of the regression lme. Compare your graph to the given one and choose the correct line. Regression line 60 50 三40 s 30 20 10 0 0 25 50 75 100 Trunk diameter (cm) The correct regression line is blue black (B) The tree diameters ranged from 1 to 101 an·Predict the height of a tree in this area that is 50 cm in diameter at breast height. (Enteryour answer rounded to two decimal places.) User to argue the reliability of this prediction. ris high implying that the prediction is very reliable, provided that the relationship is lnear ris high implying that the prediction is very reliable, provided that the relationship is non-linear r is low implying that the prediction is very reliable, provided that the relationship is non-linear ris low implying that the prediction is very reliable, provided that the relationship is limear
Answer #1

a)slope =r*Sy/sx=0.96*13.4/23.5=0.5474

intercept =Ybar-b1*xbar=15.6-0.5474*23.4=2.791

the correct regression line : Black

b)predicted height =2.791+0.5474*50=30.16 cm

r2 is high implying that the prediction is very reliable provided that the relationship is linear